An airplane flew 4 hours with a 25 mph tail wind. The return trip against the same wind took 5 hours. Find the speed of the airplane in still air.
Answer by ankor@dixie-net.com(22740)

An airplane flew 4 hours with a 25 mph tail wind. The return trip against the same wind took 5 hours. Find the speed of the airplane in still air.
:
Let s = speed in still air
:
(s+25) = speed with wind
(s-25) = speed against the wind
:
The trip out and the trip back are equal distance. Write a distance equation.
Distance = time * speed
:
5(s-25) = 4(s+25)
5s - 125 = 4s + 100
5s - 4s = 100 + 125
s = 225 mph in still air
:
:
Check the solution by finding the distances to be equal:
Against: 5(200) = 1000 mi
With: 4(250) = 1000 mi confirms our solution